OpenStudy (anonymous):

12.) The table below shows pork production in China from 2000 to 2007. Use a calculator to find the line of best fit. b. Use your linear model to predict how many metric tons of pork will be produced in 2025. c. Use your linear model to predict when production is likely to reach 100,000 metric tons. [Year: 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2007 Production(metric tons): 40,475, 42,010, 43,413, 45,331, 47,177, 50,254, 52,407, 54,491.]

OpenStudy (zarkon):

do you mean to have "010" in your data set?

OpenStudy (zarkon):

or is that 42010?...too many commas

OpenStudy (anonymous):

42,010

OpenStudy (ranga):

40475, 42010, 43413, 45331, 47177, 50254, 52407, 54491 looks better.

OpenStudy (anonymous):

I dunno what to do. i've been trying to do this all day:(

OpenStudy (ranga):

If you have a calculator that gives you the line of best fit just input the 8 data points. Choose x = 0 to represent year 2000 So the 8 data points are: (0, 40475), (1, 42010), (2, 43413), (3, 45331), (4, 47177), (5, 50254), (6, 52407), (7, 54491) You should get the slope of the line and the y intercept from the calculator. So the line will be y = mx + b. Put in x = 25 to find the pork production in 2025.

OpenStudy (anonymous):

I don't have a calc..

OpenStudy (ranga):

And for part c) set y = 100000 and solve for x.

OpenStudy (anonymous):

what do I write down on my paper?

OpenStudy (ranga):

Let me see if there is an online calculator that provides best fit equations.

OpenStudy (anonymous):

I'm so behind on this.. Errgg

OpenStudy (ranga):

Try this: http://www.alcula.com/calculators/statistics/linear-regression/ I have not used it before. But give it a shot.

OpenStudy (anonymous):

what do i do?

OpenStudy (ranga):

Well, I just plugged in the 8 x and y values and I am getting the line of best fit to be: y = 2053x + 39759 b) How much pork in 2025? Put x = 25 in the above equation: y = (2053)(25) + 39759 = 91084 So in the year 2025 the pork production will be 91084 metric tons c) When will production reach 100000 Put y = 100000 in the equation and solve for x: 100000 = 2053x + 39759 x = (100000 - 39759) / 2053 = 29.3 Round it to 29 So the pork production will reach 100000 in the year 2029.

OpenStudy (anonymous):

is that the whole answer?

OpenStudy (ranga):

In the above equation, I rounded off m and b in y = mx + b to the nearest whole number. But if I use the online calculator the exact answer to b) comes to 91088 metric tons and answer to c) comes to 2030. Answers to the question: a) Line of best fit is: y= 2053.17x + 39758.67 where x is the number of years since 2000 and y is the pork production in metric tons. b) 91088 metric tons c) 2030

OpenStudy (anonymous):

thank you